OpenStudy (anonymous):

a parabola has the given x-intercepts. what is the x coordinate of the vertex? write your result in decimal form. (-4,0) (3,0)

OpenStudy (anonymous):

Equation of the parabola: \[ y=(x+4)(x-3)=x^2+x-12 \] Vertex is at \(\left(-\frac 1 2, -\frac {49}{4}\right) \)

OpenStudy (anonymous):

Since it has the x-ints -4, 3 , that means the solution is -4, 3 Then you can form a quadratic equation with the given roots as [x-(-4)](x-3) =0 (x+4)(x-3)=0 Expand it, you'll get x^2 + 4x -3x -12 =0 x^2 +x -12 =0 Now, using the method of completing square. (x+x+ (1/2)^2) - (1/2)^2 -12 =0 (x+1/2)^2 - 49/4 =0 Since vertex is (h,k) in the equation y=a(x-h)^2+k Compare it and you will get h=-1/2, k=-49/4 So, the vertex is (-1/2 , -49/4)